A manufacturing market is initially at equilibrium at equilibrium price and quantity.suddenly the demand of its products increases by 10% while supply of products decrease by 15% . Find its new equilibrium price, equilibrium point and equilibrium quantity.


1
Expert's answer
2021-05-03T10:57:05-0400
"Q_D^\/=1.1Q_d"

"Q_S^\/=0.85Q_s"

"1.1Q_D=0.85Q_s"

"Q_D=0.77Q_S"

"a-bp=0.77(c+dp)"

"a-bp=0.77c+0.77dp"

"0.77dp+bp=a-0.77c"

"p=\frac{a-0.77c}{0.77d+b}"